# BaB-ND: Long-Horizon Motion Planning with Branch-and-Bound and Neural Dynamics

### v1

@@ B.2 Details about bound propagation early-stop
Figure A8: Bound propagation with early-stop on an $L$-layer MLP $f({\bm{u}})$. Bound propagation starts from the node of output $f$ and then backwards layer by layer to the $L-1$-th linear layer. The backward flow is highlighted as and stop at . In the queue, the node popped at every step is semitransparent.
Assume $f({\bm{u}})$ is an $L$-layer MLP. We illustrate how to estimate its lower bound $\underline{f}^{*}$ with early stopping at the last $\mathrm{ReLU}$ layer in Figure A8. For simplicity, we denote the output $f({\bm{u}})$, the $L$-th linear layer, the $\mathrm{ReLU}$ layer, the $L-1$-th linear layer, and the input ${\bm{u}}$ as nodes $\mathtt{N}_{f}$, $\mathtt{N}_{L}$, $\mathtt{N}_{R}$, $\mathtt{N}_{L-1}$, and $\mathtt{N}_{{\bm{u}}}$, respectively. Additionally, we denote $f({\bm{u}})=z^{(L)}({\bm{u}})$ as the output value of node $\mathtt{N}_{L}$, $\hat{z}^{(L-1)}({\bm{u}})$ as the input of $\mathtt{N}_{L}$ and output of $\mathtt{N}_{R}$, and $z^{(L-1)}({\bm{u}})$ as the input of $\mathtt{N}_{R}$ and output of $\mathtt{N}_{L-1}$.
@@ C.1 Scalability analysis
The results in Figure A9 show that the runtime of our BaB-ND is less sensitive to the increasing complexity of planning problems compared to sampling-based methods. While BaB-ND incurs additional overhead from initializing $\alpha,\!\beta$-CROWN and performing branching and bounding, it is less efficient than sampling-based methods for small problems.
Figure A9: Comparison of runtime with sampling-based methods. Although our BaB-ND is less efficient on small planning problems than baselines, it achieves similar efficiency on larger planning problems.
We also report the average objectives for all methods on the largest four planning problems to evaluate their effectiveness in Table A1. Overall, the performance gaps between our BaB-ND and the baselines increase with the size of the problem, highlighting the ineffectiveness of sampling-based methods for large, complex planning problems.
Table A1: Comparison of planning performance across different configurations
@@
Additionally, we evaluate the planning performance of sampling-based methods and our approach on the same simple synthetic planning problems as those in Figure 7. We report only the six cases that MIP can solve optimally within 300 seconds. The results in Table A2 show that, under these much simpler settings compared to those of our main experiments, all methods perform similarly. Sampling-based methods (MPPI, CEM, and ours) achieve a gap under the order of $1\times 10^{-4}$ compared to MIP with an optimality guarantee.
Table A2: Comparison of planning performance on simple synthetic planning problems
@@
Figure A10: Comparison of runtime on CPU and GPU.
@@
We evaluate the performance improvement from CPU to GPU in Figure A10. We use the same test cases as in Figure 7 and report “NaN” if the process does not terminate within 300 seconds.
@@ C.2 Comparison with Conventional Motion Planning Approaches
We conduct an additional experiment on the task Pushing with Obstacle to compare the planning performance of our sampling-based baselines, our BaB-ND, and two conventional motion planning approaches: 1. Rapidly-exploring Random Tree (RRT); 2. Probabilistic Roadmap (PRM). In Table A3. Since RRT and PRM do not optimize the objective as we did in sampling-based methods and our BaB-ND, we only report the step cost at planning horizon $H$ as the final step cost instead of the planning objective.
Table A3: Comparison of planning performance with RRT and PRM
@@ C.3 Ablation Study and Hyper-Parameter Analysis
We conduct an additional ablation study on the Pushing with Obstacles and Object Sorting tasks to evaluate how different design choices impact planning performance in Table A4.
Table A4: Ablation study on branching and bounding components
@@
We report the mean objectives under different hyper-parameter configurations in Table A5. The base hyper-parameter configuration is $\eta=0.75$, $T=0.05$, and $w=1$. For benchmarking, we vary at most one hyper-parameter at a time while keeping the others fixed at the base configuration.
Table A5: Planning performance under different hyper-parameter configurations
@@ C.4 Quantitative Analysis on Search Space
We conducted an experiment to measure the normalized space size of pruned subdomains over iterations. In Table A6, we report three metrics over the branch-and-bound iterations: 1. the normalized space size of pruned subdomains, 2. the size of the selected subdomains, and 3. the improvement in the objective value.
@@
Table A6: Performance Metrics Over Iterations
@@ C.5 Performance Change with Varying Input Discontinuities
We conducted a follow-up experiment by removing the obstacles (non-feasible regions) in the problem of Pushing with Obstacles, simplifying the objective function. Below, we report the performance of different methods on the simplified objective function (w/o obstacles) and the original objective function (w/ obstacles) in Table A7.
@@
Table A7: Performance comparison varying input discontinuities
@@
(b) Objective w/ obstacles (Table A1)
@@ C.6 Further Scalability Analysis on the Synthetic Example
The results in Table A8 show that our BaB-ND much outperforms baselines when the input dimension increases. These results are expected since existing sampling-based methods search for solutions across the entire input space, requiring an exponentially increasing number of samples to achieve sufficient coverage. In contrast, our BaB-ND strategically splits and prunes unpromising regions of the input space, guiding and improving the effectiveness of existing sampling-based methods.
Table A8: Performance comparison across different input dimensions (Metric: Gap to $f^{*}$, $\downarrow$)
@@
We further report the following metrics about our BaB-ND in Table A9 to better understand the behavior of BaB-ND under high-dimensional cases: 1. The gap between the best objective found and the optimal objective value as above, 2. The normalized space size of pruned subdomains at the last iteration, 3. The normalized space size of selected subdomains at the last iteration, and 4. The total runtime.
@@
Table A9: Performance metrics across different input dimensions $N$
@@ D.1 Definition of Actions and States
The action is defined as a 2D movement of the pusher, $(\Delta x,\Delta y)$, within a single step, as illustrated in Figure A11.a. Four keypoints are assigned to the “T”-shaped object to capture its geometric features, as visualized in Figure A12.a.
@@
The action is similarly defined as a 2D movement of the pusher, $(\Delta x,\Delta y)$, within one step, as shown in Figure A11.b. The orientation of the pusher rotates around the gravity direction and is always perpendicular to the direction of pushing $(\Delta x,\Delta y)$. The state is represented by six keypoints on the “L”-shaped objects (three keypoints per object), as depicted in Figure A12.b.
@@
This task involves 3 DoF actions, defined as movements of the robot gripper along the Cartesian axes, $(\Delta x,\Delta y,\Delta z)$, as illustrated in Figure A11.c. To better represent the rope’s geometric configuration, multiple keypoints (e.g., 10) are sampled along the rope, as shown in Figure A12.c.
@@
This task requires a series of independent long-range pushing actions, where the initial position of the pusher for each step is independent of its final position from the previous step. The action is defined by the 2D initial position of the pusher, $(x,y)$, and its subsequent 2D movement, $(\Delta x,\Delta y)$, as visualized in Figure A11.d. For the state, we use an object-centric representation, defining the state as the center points of all objects, as illustrated in Figure A12.d. Interactions between objects are modeled as a graph using graph neural networks.
@@
Figure A11: Visualization of real-world experiments with illustration of action spaces in different tasks. In tasks Pushing w/ Obstacles and Object Merging, the actions are defined as the movement of the pusher in 2D space. In the task Rope Routing, the action is defined as the movement of the gripper in 3D space. In the task Object Sorting, the 4 DoF action is defined as the 2D starting position and 2D movement of the pusher.
@@ D.2 Data Collection
As shown in Figure A12.a, the pusher is simulated as a 5mm cylinder. The stem of the “T”-shaped object has a length of 90mm and a width of 30mm, while the bar has a length of 120mm and a width of 30mm. The pushing action along the x-y axis is limited to 30mm. We don’t add explicit obstacles in the data generation process, while the obstacles are added as penalty terms during planning. We generated 32,000 episodes, each containing 30 pushing actions between the pusher and the “T”-shaped object.
@@
As shown in Figure A12.b, the pusher is simulated as a 5mm cylinder. The leg of the “L”-shaped object has a length of 30mm and a width of 30mm, while the foot has a length of 90mm and a width of 30mm. The pushing action along the x-y axis is limited to 30mm. We generated 64,000 episodes, each containing 40 pushing actions between the pusher and the two “L”-shaped objects.
@@
As shown in Figure A12.c, we use an xArm6 robot with a gripper to interact with the rope. The rope has a length of 30cm and a radius of 0.03cm. One end of the rope is fixed while the gripper grasps the other end. We randomly sample actions in 3D space, with the action bound set to 30cm. The constraint is that the distance between the gripper position and the fixed end of the rope cannot exceed the rope length. We generated 15,000 episodes, each containing 6 random actions. For this task, we post-process the dataset and split each action into 2cm sections.
@@
As shown in Figure A12.d, the pusher is simulated as a rectangle measuring 45mm by 3.5mm. The radius of the object pieces is set to 15mm. For this task, we use long push as our action representation, which generates the start position and pushing action length along the x-y axis. The pushing action length is bounded between -100mm and 100mm. We generated 32,000 episodes, each containing 12 pushing actions between the pusher and the object pieces.
Figure A12: Simulation environments used for data collection. We use Pymunk to simulate environments involving only rigid body interactions. For manipulating the deformable rope, we utilize NVIDIA FleX to simulate the interactions between the rope and the robot gripper.
@@ D.6 Hyperparameters of Baselines
We follow previous works (Li et al., 2018; 2019) to implement gradient descent (GD) for planning and enhance it through hyperparameter tuning of the step size. Specifically, we define a base step size and multiply it by varying ratios to launch independent instances with different step sizes. The best result among all instances is recorded for performance comparison. Similarly, we adapt MPPI (Williams et al., 2017) by tuning its noise level and reward temperature. We also implement the Decentralized Cross-Entropy Method (CEM) (Zhang et al., 2022c), which enhances CEM using an ensemble of independent instances. The specific hyperparameter settings for these methods are provided in Table A10.
Table A10: Hyperparameter setting for baselines

### v2

@@ B.2 Details about bound propagation early-stop
Figure 8: Bound propagation with early-stop on an $L$-layer MLP $f({\bm{u}})$. Bound propagation starts from the node of output $f$ and then backwards layer by layer to the $L-1$-th linear layer. The backward flow is highlighted as and stop at . In the queue, the node popped at every step is semitransparent.
Assume $f({\bm{u}})$ is an $L$-layer MLP. We illustrate how to estimate its lower bound $\underline{f}^{*}$ with early stopping at the last $\mathrm{ReLU}$ layer in Figure 8. For simplicity, we denote the output $f({\bm{u}})$, the $L$-th linear layer, the $\mathrm{ReLU}$ layer, the $L-1$-th linear layer, and the input ${\bm{u}}$ as nodes $\mathtt{N}_{f}$, $\mathtt{N}_{L}$, $\mathtt{N}_{R}$, $\mathtt{N}_{L-1}$, and $\mathtt{N}_{{\bm{u}}}$, respectively. Additionally, we denote $f({\bm{u}})=z^{(L)}({\bm{u}})$ as the output value of node $\mathtt{N}_{L}$, $\hat{z}^{(L-1)}({\bm{u}})$ as the input of $\mathtt{N}_{L}$ and output of $\mathtt{N}_{R}$, and $z^{(L-1)}({\bm{u}})$ as the input of $\mathtt{N}_{R}$ and output of $\mathtt{N}_{L-1}$.
@@ C.1 Scalability analysis
The results in Figure 9 show that the runtime of our BaB-ND is less sensitive to the increasing complexity of planning problems compared to sampling-based methods. While BaB-ND incurs additional overhead from initializing $\alpha,\!\beta$-CROWN and performing branching and bounding, it is less efficient than sampling-based methods for small problems.
Figure 9: Comparison of runtime with sampling-based methods. Although our BaB-ND is less efficient on small planning problems than baselines, it achieves similar efficiency on larger planning problems.
We also report the average objectives for all methods on the largest four planning problems to evaluate their effectiveness in Table 1. Overall, the performance gaps between our BaB-ND and the baselines increase with the size of the problem, highlighting the ineffectiveness of sampling-based methods for large, complex planning problems.
Table 1: Comparison of planning performance across different configurations
@@
Additionally, we evaluate the planning performance of sampling-based methods and our approach on the same simple synthetic planning problems as those in Figure 7. We report only the six cases that MIP can solve optimally within 300 seconds. The results in Table 2 show that, under these much simpler settings compared to those of our main experiments, all methods perform similarly. Sampling-based methods (MPPI, CEM, and ours) achieve a gap under the order of $1\times 10^{-4}$ compared to MIP with an optimality guarantee.
Table 2: Comparison of planning performance on simple synthetic planning problems
@@
Figure 10: Comparison of runtime on CPU and GPU.
@@
We evaluate the performance improvement from CPU to GPU in Figure 10. We use the same test cases as in Figure 7 and report “NaN” if the process does not terminate within 300 seconds.
@@ C.2 Comparison with Conventional Motion Planning Approaches
We conduct an additional experiment on the task Pushing with Obstacle to compare the planning performance of our sampling-based baselines, our BaB-ND, and two conventional motion planning approaches: 1. Rapidly-exploring Random Tree (RRT); 2. Probabilistic Roadmap (PRM). In Table 3. Since RRT and PRM do not optimize the objective as we did in sampling-based methods and our BaB-ND, we only report the step cost at planning horizon $H$ as the final step cost instead of the planning objective.
Table 3: Comparison of planning performance with RRT and PRM
@@ C.3 Ablation Study and Hyper-Parameter Analysis
We conduct an additional ablation study on the Pushing with Obstacles and Object Sorting tasks to evaluate how different design choices impact planning performance in Table 4.
Table 4: Ablation study on branching and bounding components
@@
We report the mean objectives under different hyper-parameter configurations in Table 5. The base hyper-parameter configuration is $\eta=0.75$, $T=0.05$, and $w=1$. For benchmarking, we vary at most one hyper-parameter at a time while keeping the others fixed at the base configuration.
Table 5: Planning performance under different hyper-parameter configurations
@@ C.4 Quantitative Analysis on Search Space
We conducted an experiment to measure the normalized space size of pruned subdomains over iterations. In Table 6, we report three metrics over the branch-and-bound iterations: 1. the normalized space size of pruned subdomains, 2. the size of the selected subdomains, and 3. the improvement in the objective value.
@@
Table 6: Performance Metrics Over Iterations
@@ C.5 Performance Change with Varying Input Discontinuities
We conducted a follow-up experiment by removing the obstacles (non-feasible regions) in the problem of Pushing with Obstacles, simplifying the objective function. Below, we report the performance of different methods on the simplified objective function (w/o obstacles) and the original objective function (w/ obstacles) in Table 7.
@@
Table 7: Performance comparison varying input discontinuities
@@
(b) Objective w/ obstacles (Table 1)
@@ C.6 Further Scalability Analysis on the Synthetic Example
The results in Table 8 show that our BaB-ND much outperforms baselines when the input dimension increases. These results are expected since existing sampling-based methods search for solutions across the entire input space, requiring an exponentially increasing number of samples to achieve sufficient coverage. In contrast, our BaB-ND strategically splits and prunes unpromising regions of the input space, guiding and improving the effectiveness of existing sampling-based methods.
Table 8: Performance comparison across different input dimensions (Metric: Gap to $f^{*}$, $\downarrow$)
@@
We further report the following metrics about our BaB-ND in Table 9 to better understand the behavior of BaB-ND under high-dimensional cases: 1. The gap between the best objective found and the optimal objective value as above, 2. The normalized space size of pruned subdomains at the last iteration, 3. The normalized space size of selected subdomains at the last iteration, and 4. The total runtime.
@@
Table 9: Performance metrics across different input dimensions $N$
@@ D.1 Definition of Actions and States
The action is defined as a 2D movement of the pusher, $(\Delta x,\Delta y)$, within a single step, as illustrated in Figure 11.a. Four keypoints are assigned to the “T”-shaped object to capture its geometric features, as visualized in Figure 12.a.
@@
The action is similarly defined as a 2D movement of the pusher, $(\Delta x,\Delta y)$, within one step, as shown in Figure 11.b. The orientation of the pusher rotates around the gravity direction and is always perpendicular to the direction of pushing $(\Delta x,\Delta y)$. The state is represented by six keypoints on the “L”-shaped objects (three keypoints per object), as depicted in Figure 12.b.
@@
This task involves 3 DoF actions, defined as movements of the robot gripper along the Cartesian axes, $(\Delta x,\Delta y,\Delta z)$, as illustrated in Figure 11.c. To better represent the rope’s geometric configuration, multiple keypoints (e.g., 10) are sampled along the rope, as shown in Figure 12.c.
@@
This task requires a series of independent long-range pushing actions, where the initial position of the pusher for each step is independent of its final position from the previous step. The action is defined by the 2D initial position of the pusher, $(x,y)$, and its subsequent 2D movement, $(\Delta x,\Delta y)$, as visualized in Figure 11.d. For the state, we use an object-centric representation, defining the state as the center points of all objects, as illustrated in Figure 12.d. Interactions between objects are modeled as a graph using graph neural networks.
@@
Figure 11: Visualization of real-world experiments with illustration of action spaces in different tasks. In tasks Pushing w/ Obstacles and Object Merging, the actions are defined as the movement of the pusher in 2D space. In the task Rope Routing, the action is defined as the movement of the gripper in 3D space. In the task Object Sorting, the 4 DoF action is defined as the 2D starting position and 2D movement of the pusher.
@@ D.2 Data Collection
As shown in Figure 12.a, the pusher is simulated as a 5mm cylinder. The stem of the “T”-shaped object has a length of 90mm and a width of 30mm, while the bar has a length of 120mm and a width of 30mm. The pushing action along the x-y axis is limited to 30mm. We don’t add explicit obstacles in the data generation process, while the obstacles are added as penalty terms during planning. We generated 32,000 episodes, each containing 30 pushing actions between the pusher and the “T”-shaped object.
@@
As shown in Figure 12.b, the pusher is simulated as a 5mm cylinder. The leg of the “L”-shaped object has a length of 30mm and a width of 30mm, while the foot has a length of 90mm and a width of 30mm. The pushing action along the x-y axis is limited to 30mm. We generated 64,000 episodes, each containing 40 pushing actions between the pusher and the two “L”-shaped objects.
@@
As shown in Figure 12.c, we use an xArm6 robot with a gripper to interact with the rope. The rope has a length of 30cm and a radius of 0.03cm. One end of the rope is fixed while the gripper grasps the other end. We randomly sample actions in 3D space, with the action bound set to 30cm. The constraint is that the distance between the gripper position and the fixed end of the rope cannot exceed the rope length. We generated 15,000 episodes, each containing 6 random actions. For this task, we post-process the dataset and split each action into 2cm sections.
@@
As shown in Figure 12.d, the pusher is simulated as a rectangle measuring 45mm by 3.5mm. The radius of the object pieces is set to 15mm. For this task, we use long push as our action representation, which generates the start position and pushing action length along the x-y axis. The pushing action length is bounded between -100mm and 100mm. We generated 32,000 episodes, each containing 12 pushing actions between the pusher and the object pieces.
Figure 12: Simulation environments used for data collection. We use Pymunk to simulate environments involving only rigid body interactions. For manipulating the deformable rope, we utilize NVIDIA FleX to simulate the interactions between the rope and the robot gripper.
@@ D.6 Hyperparameters of Baselines
We follow previous works (Li et al., 2018; 2019) to implement gradient descent (GD) for planning and enhance it through hyperparameter tuning of the step size. Specifically, we define a base step size and multiply it by varying ratios to launch independent instances with different step sizes. The best result among all instances is recorded for performance comparison. Similarly, we adapt MPPI (Williams et al., 2017) by tuning its noise level and reward temperature. We also implement the Decentralized Cross-Entropy Method (CEM) (Zhang et al., 2022c), which enhances CEM using an ensemble of independent instances. The specific hyperparameter settings for these methods are provided in Table 10.
Table 10: Hyperparameter setting for baselines